# Born a Transformer – Always a Transformer?

### v2

@@ A.2 Expressiveness Results
C-Rasp[pos] program for UR, UL $PRED_{a}(i)$ $:=\exists[j\leq i,j=i-1]{Q_{a}(j)}$ (1) for each $a\in\Sigma$ $CBIGRAM_{ab}$ $:=\exists[j\leq i]{Q_{b}(j)\land PRED_{a}(j)}$ (2) for each $a,b\in\Sigma$ $NEXT_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land CBIGRAM_{\sigma a}(i)\right]$ (UR version) (3) for each $a\in\Sigma$ $NEXT_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land CBIGRAM_{a\sigma}(i)\right]$ (UL version) (4) for each $a\in\Sigma$
This program is a direct formalization of the (anti-)induction head circuit, and basic to the other C-Rasp[pos] programs constructed below. The first line checks for each $a\in\Sigma$, at each position $i$, whether position $i-1$ holds the symbol $a$. The second line checks, for each bigram $ab\in\Sigma\times\Sigma$ whether it appears at some position in the context up to the $i$-th position. The third line then states that, in UR, the correct next token is $a$ if and only if the bigram $\sigma a$ has appeared in the context, where $\sigma$ is the symbol at position $i$. The fourth line is the analogous instruction for UL: Here, $a$ is predicted if and only if the bigram $a\sigma$ appears in the context. ∎
@@
C-Rasp[pos] program for NRFirst $ISLEFTMOST(i)$ $:=\bigvee_{a\in\Sigma}\left[Q_{a}(i)\wedge\left(\#[j\leq i]{Q_{a}(j)}\right)\leq 1\right]$ (1) $PRED_{a}(i)$ $:=\exists[j\leq i,j=i-1]\left[Q_{a}(j)\wedge ISLEFTMOST(j)\right]$ (2) for each $a\in\Sigma$ $CBIGRAM_{ab}$ $:=\exists[j\leq i]\left[{Q_{b}(j)\land PRED_{a}(j)}\right]$ (3) for each $a,b\in\Sigma$ $NEXT_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land CBIGRAM_{\sigma a}(i)\right]$ (4) for each $a\in\Sigma$
where $NEXT_{a}(i)$ holds at the final position if and only if the desired completion is the symbol $a$. The first line in the program checks, at each position, if it is the leftmost representative of the symbol that it holds: That is, whether there is a symbol $a\in\Sigma$ that occurs at position $i$ but at no earlier position. For each $a\in\Sigma$, the second line checks, at position $i$, whether the preceding position $i-1$ holds the symbol $a$ – simulating the operation of a Previous-Token head. Additionally it also uses the computations made in the first line to check whether position $i-1$ was the leftmost representative with such a property. For each possible bigram $ab\in\Sigma\times\Sigma$, the third line checks whether it has appeared or not. Finally, the fourth line says that $a$ is the correct next token if and only if the current position holds a symbol $\sigma$ such that $\sigma a$ appears as a bigram in the context. This last line is the one that simulates the operation of the Induction head itself.
@@
C-Rasp[pos] program for NLFirst $ISLEFTMOST(i)$ $:=\bigvee_{a\in\Sigma}\left[Q_{a}(i)\wedge\neg\exists[j\leq i]{Q_{a}(j)}\right]$ (1) $PRED_{a}(i)$ $:=\exists[j\leq i,j=i-1]\left[Q_{a}(j)\wedge ISLEFTMOST_{a}(j)\right]$ (2) for each $a\in\Sigma$ $CBIGRAM_{ab}$ $:=\exists[j\leq i]\left[{Q_{b}(j)\land PRED_{a}(j)}\right]$ (3) for each $a,b\in\Sigma$ $NEXT_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{a}(i)\land CBIGRAM_{a\sigma}(i)\right]$ (4) for each $a\in\Sigma$
where $NEXT_{a}(i)$ holds at the final position if and only if the desired completion is the symbol $a$.
@@
We show this by reducing to Lemma 36 in Huang et al. [2025]. Assume, for the sake of contradiction, that NRLast was expressible in C-Rasp[pos]. This program would discriminate between the two languages
[…]
@@ A.2.1 Generalized reverse copying
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\rho_{\mathrm{word}}$ reverses a list of chunks, whereas $\rho_{\mathrm{tok}}$ does the same when every chunk is a single token. When each word has length $1$ the two coincide, so solving word-order reversal would solve Non-unique copy-backwards as a special case—which creates a contradiction, and thus $\rho_{\mathrm{word}}$ should also not be solvable.
@@ Appendix B From Scratch Training Experiments
We train models with the default CausalLMLoss from the transformers library, the length of inputs is sampled uniformly from minimum up to maximum length in the specified range. This is true for training data where the range is $(l_{min},50)$ where $l_{min}$ is the minimum length possible for a given task (varies from 2 - 4, depending on the task variant and the minimum length of characters required to make the string valid). Our test sets have the size – $(l_{min},50),(51,100),(101,150)$. At each step, the model outputs the correct continuation – for retrieval the correct token to be retrieved, for copying the subsequent next token from the input sequence. The models are trained on a whole sequence of tokens. We train decoder-only transformer from scratch, using implementations from HuggingFace transformers for APE (GPT2LMHeadModel) and RoPE (LlamaForCausalLM). We train models for maximum $30$K steps with a batch size of $64$. We stop training early once the model’s accuracy reaches 100% on the in-distribution test set (the one in range $[l_{min},50]$. The model is trained with a dropout rate of $0.0$, and we use AdamW, with a weight decay rate of $0.01$ (choices based on Huang et al. [2025]).
For experiments with APE, at training time, we add random offsets to position indices so that all position embeddings are trained (following Huang et al. [2025]). The offsets are sampled uniformly at random in the range $[0,150-\ell_{curr}]$ where $\ell_{curr}$ is the length of the current string. For RoPE, such an offset is not required in training [^9] In fact, relative positional encodings such as RoPE are by definition invariant to the addition of a constant offset to the indices; instead a scaling factor of $32.0$ is set; the RoPE scaling type used was linear.
In preliminary experiments, we found that different model architectures, while achieving $100\%$ accuracy on in-distribution data, may perform differently on out-of-distribution data. To draw a conclusion about how the model performs on a problem in general, we determine the hyperparameters as follows: We consider configurations of {1, 2, 4} layers, {1, 2, 4} heads and model dimension of {16, 64, 256}, and learning rate of {0.001, 0.0001}. We sweep all the configurations by iterating over every combination and choose the one that achieves the highest accuracy on $[51,100]$ among those configurations whose accuracy on $[l_{min},50]$ is 100%. When there are multiple such options, e.g., their accuracy on $[51,100]$ is 100%, the one with the simplest architecture is selected (when estimating complexity, we assume the following priority: number of layers $>$ number of heads $>$ model dimension). The final hyperparameters we used for each task are shown in Table 1 and 2. After we determine the hyperparameter configuration, we run the experiments with multiple random seeds and report the average accuracy of 3 successful runs (those runs where the model achieves 100% accuracy on in-distribution data).
@@
Appendix C Prompting Details
C.1 Prompting: In-Context Retrieval Templates
@@
Few-shot small, same
@@
Few-Shot Pools
  • small — Examples drawn from a held-out pool with lengths smaller than the input string $X$.
@@
C.1.1 Prompting Retrieval Detailed Results and Graphs
Figure 10: In context accuracy for Qwen2.5-7B and Llama-3 8B across all our tasks averaged over 3 seeds on the best prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 11: In context accuracy for Qwen2.5-7B-Instruct and Llama-3-8B-Instruct across all our tasks averaged over 3 seeds on the best prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 12: In context accuracy for Qwen2.5-32B-Instruct and Llama-3 70B-Instruct across all our tasks averaged over 3 seeds on the best prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
As evident from all the figures attached here (smaller completion models - Figure 10), (smaller instruct models - Figure 11), (bigger instruct models - Figure 12), the left right asymmetry persists across models sizes and types. While in instruction tuned models and with bigger models, the left and the backward accuracy starts to rise, there is still always a pronounced gap within every task variant.
Amongst the prompts there was high variability for all our retrieval setups, but close to none for the copying setups. Our bare template performed the worst in most retrieval setups, and the highest accuracy metrics were achieved when the instruction template was – MATH_RULE_EXPLAINED, with the few shot setting as SMALL with the presence of a separator. For copying, just the BARE setup was enough to get high accuracy numbers without a need to specify any instructions (especially for the forward copying cases). Providing instructions did improve the performance especially for the Qwen Instruct models.
C.2 Prompting: In-Context Copying Templates
@@
For our In context Copying task suite, we use three lightweight prompt styles. There is no Cartesian grid as in retrieval, as each of the prompts had statistically little variance, and eliciting this ability from our LLMs was not as challenging as in retrieval.
  • bare — few shot examples only, no instructional text.
  • obey — a one-line English rule (rule_simple) introducing the few-shot block.
  • hint - a short hint (rule_hint) that highlights the input–output relation; mainly useful for backward copying.
@@
The phrases referenced above are drawn from Table 6. Note that unique and non-unique regimes share identical wording because the operation (copy vs. reverse) is independent of token repetition once the input is fixed.
@@
The corresponding rule_hint strings swap the lead-in "The output is …" with "In every example the output …" to make the instruction less imperative.
Few-shot sampler.
All copy prompts use the same held-out $k{=}5$-shot pool (length $L{=}5$, $N{=}1500$) that is disjoint from the evaluation set. Each prompt is built by picking five random lines from this pool, followed by the query built from the current test string.
End-to-end prompt bare template example (UF)
@@
C.3 Promping: Lorem Ipsum Copying

### v3

@@ A.2 Expressiveness Results
C-Rasp[pos] program for UR, UL $\text{PRED}_{a}(i)$ $:=\exists[j\leq i,j=i-1]{Q_{a}(j)}$ (1) for each $a\in\Sigma$ $CBIGRAM_{ab}$ $:=\exists[j\leq i]{Q_{b}(j)\land\text{PRED}_{a}(j)}$ (2) for each $a,b\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land\text{CBIGRAM}_{\sigma a}(i)\right]$ (UR version) (3) for each $a\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land\text{CBIGRAM}_{a\sigma}(i)\right]$ (UL version) (4) for each $a\in\Sigma$
This program is a direct formalization of the (anti-)induction head circuit, and basic to the other C-Rasp[pos] programs constructed below. The first line checks for each $a\in\Sigma$, at each position $i$, whether position i1i-1 holds the symbol $a$. The second line checks, for each bigram $ab\in\Sigma\times\Sigma$ whether it appears at some position in the context up to the $i$-th position. The third line then states that, in UR, the correct next token is $a$ if and only if the bigram $\sigma a$ has appeared in the context, where $\sigma$ is the symbol at position $i$. The fourth line is the analogous instruction for UL: Here, $a$ is predicted if and only if the bigram $a\sigma$ appears in the context. ∎
@@
C-Rasp[pos] program for NRFirst $\text{ISLEFTMOST}(i)$ $:=\bigvee_{a\in\Sigma}\left[Q_{a}(i)\wedge\left(\#[j\leq i]{Q_{a}(j)}\right)\leq 1\right]$ (1) $\text{PRED}_{a}(i)$ $:=\exists[j\leq i,j=i-1]\left[Q_{a}(j)\wedge\text{ISLEFTMOST}(j)\right]$ (2) for each $a\in\Sigma$ $\text{CBIGRAM}_{ab}$ $:=\exists[j\leq i]\left[{Q_{b}(j)\land\text{PRED}_{a}(j)}\right]$ (3) for each $a,b\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{\sigma}(i)\land\text{CBIGRAM}_{\sigma a}(i)\right]$ (4) for each $a\in\Sigma$
where $\text{NEXT}_{a}(i)$ holds at the final position if and only if the desired completion is the symbol $a$. The first line in the program checks, at each position, if it is the leftmost representative of the symbol that it holds: That is, whether there is a symbol $a\in\Sigma$ that occurs at position $i$ but at no earlier position. For each $a\in\Sigma$, the second line checks, at position $i$, whether the preceding position i1i-1 holds the symbol $a$ – simulating the operation of a Previous-Token head. Additionally it also uses the computations made in the first line to check whether position i1i-1 was the leftmost representative with such a property. For each possible bigram $ab\in\Sigma\times\Sigma$, the third line checks whether it has appeared or not. Finally, the fourth line says that $a$ is the correct next token if and only if the current position holds a symbol $\sigma$ such that $\sigma a$ appears as a bigram in the context. This last line is the one that simulates the operation of the Induction head itself.
@@
C-Rasp[pos] program for NLFirst $\text{ISLEFTMOST}(i)$ $:=\bigvee_{a\in\Sigma}\left[Q_{a}(i)\wedge\neg\exists[j\leq i]{Q_{a}(j)}\right]$ (1) $\text{PRED}_{a}(i)$ $:=\exists[j\leq i,j=i-1]\left[Q_{a}(j)\wedge\text{ISLEFTMOST}_{a}(j)\right]$ (2) for each $a\in\Sigma$ $\text{CBIGRAM}_{ab}$ $:=\exists[j\leq i]\left[{Q_{b}(j)\land\text{PRED}_{a}(j)}\right]$ (3) for each $a,b\in\Sigma$ $\text{NEXT}_{a}(i)$ $:=\bigvee_{\sigma\in\Sigma}\left[Q_{a}(i)\land\text{CBIGRAM}_{a\sigma}(i)\right]$ (4) for each $a\in\Sigma$
where $\text{NEXT}_{a}(i)$ holds at the final position if and only if the desired completion is the symbol $a$.
@@
We show this by reducing them to Lemma 36 in Huang et al. [2025]. Assume, for the sake of contradiction, that NRLast was expressible in C-Rasp[pos]. This program would discriminate between the two languages
$(w0|w1|i)^{*}w1i^{*}\langle sep\rangle w\text{\ \ \ \ \ \ \ \ \ \ \ \ (retrieving a ``1'')}$ (5)
@@ A.2.1 Generalized reverse copying
Let the token alphabet be $\Sigma=\{\sigma_{1},\dots,\sigma_{z}\}\cup\{\mathord{\square}\}$ where $\mathord{\square}$ represents separators. $\Sigma_{\mathrm{w}}=\Sigma\setminus\{\mathord{\square}\}$ are all tokens that can be part of a word. Define the language $L=(\Sigma_{\mathrm{w}}^{+}\mathord{\square})^{*}\Sigma_{\mathrm{w}}^{+}.$ Every $x\in L$ can be written uniquely as $x=w_{1}\mathord{\square}w_{2}\mathord{\square}\cdots\mathord{\square}w_{n}$ with $w_{i}\in\Sigma_{\mathrm{w}}^{+}$ (i.e. each $w_{i}$ represents a word.) and also as $x=t_{1}\dots t_{m}$, where each $t_{i}\in\Sigma$ and is the token level representation of the string. We can think of copying this string $x\in L$ in 2 ways.
@@
$\rho_{\mathrm{word}}$ is not expressible in $\mathsf{C\text{-}\textsc{Rasp}[pos]}$.
@@
Theorem 5 states that Non‑Unique Copy backwards which is the same as $\rho_{\mathrm{tok}}$—lies outside $\mathsf{C\text{-}\textsc{Rasp}[pos]}$ by a communication‑complexity argument. To transfer this inexpressibility to $\rho_{\mathrm{word}}$ consider the following reduction.
@@
and deleting the $\mathord{\square}$ symbols recovers $\rho_{\mathrm{tok}}(t_{1}\ldots t_{m})$. Hence an algorithm for $\rho_{\mathrm{word}}$ would give one for $\rho_{\mathrm{tok}}$ with constant overhead, contradicting Theorem 5. ∎
@@
$\rho_{\mathrm{word}}$ reverses a list of chunks, whereas $\rho_{\mathrm{tok}}$ does the same when every chunk is a single token. When each word has length $1$ the two coincide, so solving word‑order reversal would solve Non-unique copy‑backwards as a special case—which creates a contradiction, and thus $\rho_{\mathrm{word}}$ should also not be solvable.
@@ Appendix B From Scratch Training Experiments
We train models with the default CausalLMLoss from the transformers library, the length of inputs is sampled uniformly from minimum up to maximum length in the specified range. This is true for training data where the range is $(l_{min},50)$ where $l_{min}$ is the minimum length possible for a given task (varies from 2 - 4, depending on the task variant and the minimum length of characters required to make the string valid). Our test sets have the size – $(l_{min},50),(51,100),(101,150)$. At each step, the model outputs the correct continuation – for retrieval the correct token to be retrieved, for copying the subsequent next token from the input sequence. The models are trained on a whole sequence of tokens. We train decoder-only transformer from scratch, using implementations from HuggingFace transformers for APE (GPT2LMHeadModel) and RoPE (LlamaForCausalLM). We train models for maximum 3030K steps with a batch size of $64$. We stop training early once the model’s accuracy reaches 100% on the in-distribution test set (the one in range $l_{min},50]$. The model is trained with a dropout rate of $0.0$, and we use AdamW, with a weight decay rate of $0.01$ (choices based on Huang et al. [2025]).
For experiments with APE, at training time, we add random offsets to position indices so that all position embeddings are trained (following Huang et al. [2025]). The offsets are sampled uniformly at random in the range $0,150-\ell_{curr}]$ where $\ell_{curr}$ is the length of the current string. For RoPE, such an offset is not required in training [^10] In fact, relative positional encodings such as RoPE are by definition invariant to the addition of a constant offset to the indices; instead a scaling factor of $32.0$ is set; the RoPE scaling type used was linear.
In preliminary experiments, we found that different model architectures, while achieving $100\%$ accuracy on in-distribution data, may perform differently on out-of-distribution data. To draw a conclusion about how the model performs on a problem in general, we determine the hyperparameters as follows: We consider configurations of {1, 2, 4} layers, {1, 2, 4} heads and model dimension of {16, 64, 256}, and learning rate of {0.001, 0.0001}. We sweep all the configurations by iterating over every combination and choose the one that achieves the highest accuracy on $[51,100]$ among those configurations whose accuracy on $l_{min},50]$ is 100%. When there are multiple such options, e.g., their accuracy on $[51,100]$ is 100%, the one with the simplest architecture is selected (when estimating complexity, we assume the following priority: number of layers >> number of heads >> model dimension). The final hyperparameters we used for each task are shown in Table 1 and 2. After we determine the hyperparameter configuration, we run the experiments with multiple random seeds and report the average accuracy of 3 successful runs (those runs where the model achieves 100% accuracy on in-distribution data).
@@
Appendix C Additional experimental details and results
C.1 Prompting: In-Context Retrieval Templates (Section 4.1)
@@ Appendix C Additional experimental details and results
Few‑shot small, same
@@
Few‑Shot Pools
  • small — Examples drawn from a held‑out pool with lengths smaller than the input string $X$.
@@
C.2 Prompting: In-Context Copying Templates (Section 4.1)
@@
For our In-context Copying task suite, we use three lightweight prompt styles. There is no Cartesian grid as in retrieval, as each of the prompts had statistically little variance, and eliciting this ability from our LLMs was not as challenging as in retrieval.
  • bare — few-shot examples only, no instructional text.
  • obey — a one‑line English rule (rule_simple) introducing the few‑shot block.
  • hint — a short hint (rule_hint) that highlights the input–output relation; mainly useful for backward copying.
@@
The phrases referenced above are drawn from Table 6. Note that unique and non‑unique regimes share identical wording because the operation (copy vs. reverse) is independent of token repetition once the input is fixed.
@@
The corresponding rule_hint strings swap the lead‑in "The output is …" with "In every example the output …" to make the instruction less imperative.
Few‑shot sampler.
All copy prompts use the same held‑out k=5k{=}5‑shot pool (length L=5L{=}5, N=1500N{=}1500) that is disjoint from the evaluation set. Each prompt is built by picking five random lines from this pool, followed by the query built from the current test string.
End‑to‑end prompt bare template example (UF)
@@
C.3 Longer Vocabulary (Word-Level) Experiments (Section 4.1)
To ensure that the Uniqueness and Directional Biases observed in our primary character-level experiments are not simply artifacts of a synthetic setup, we conducted a parallel set of experiments using a more naturalistic word-level vocabulary. In case of LLMs, the distinction between a token-level and a word-level task may be less sharp than it appears. Kaplan et al. [2025] shows that LLMs form an “inner lexicon” where simple or common words are processed as cohesive semantic units. Given their results, this word-level task formulation can be considered analogous to the character-level task.
Experimental Setup
The setup for the word-level vocabulary mirrors the main experiments in Section 4.1 but replaces the character-based vocabulary with a vocabulary of 300+ unique English words. This allowed us to construct longer, more complex sequences with lengths up to 300 tokens. All other aspects of the methodology, few-shot prompting strategy, and task structures (e.g. UR vs. UL, UF vs. UB) remain identical. The exact vocabulary we used to generate the test strings in all of the experiments is as follows.
Character-level vocabulary
a, à, â, ā, ä, ą, b, c, ç, ć, ĉ, ċ, č, d, e, é,
ê, ë, ě, ē, ę, f, g, ĝ, ğ, h, ĥ, i, í, j, ĵ, k,
l, m, n, o, ó, ô, ö, õ, p, q, r, s, t, u, ú, ù,
û, ü, v, w, x, y, z,
A, Á, Â, Ã, Ä, Å, Æ, B, C, Ç, Ć, Ĉ, Ċ, Č, D, E,
È, É, Ē, Ė, F, G, H, I, Í, Ì, Î, Ï, Ī, Ĭ, Ĩ, J,
K, L, Ĺ, M, N, Ń, O, Ó, Ô, Ö, P, Q, R, S, T, U,
Ú, Ù, Û, Ü, V, W, X, Y, Z,
0, 1, 2, 3, 4, 5, 6, 7, 8, 9
Word-level vocabulary
apple, ant, arrow, anchor, artist, animal, angle, apricot, arch, armor, axis, avenue
ball, bat, book, bridge, bottle, bucket, bench, bread, bell, button, brush, branch
cat, car, cup, cloud, clock, candle, coin, chair, circle, crown, castle, cookie
dog, door, desk, drum, duck, doll, diamond, dish, dress, dream, drop, dust
egg, ear, eye, engine, elbow, earth, envelope, exit, echo, edge, event, energy
fish, fan, fork, flower, flag, feather, fire, frame, forest, farm, fruit, fence
goat, game, glass, glove, gate, garden, gift, grape, guitar, gold, gear, group
hat, hand, horse, house, hill, hammer, heart, honey, hook, horn, hug, hope
ice, iron, ink, island, idea, image, item, ivory, icon, input, issue, idol
jar, jam, jet, jewel, jungle, jacket, juice, job, joke, joy, judge, jump
kite, key, king, knee, kitchen, knife, kitten, knight, kick, kettle, kind, koala
lamp, leaf, lion, lock, ladder, lake, lemon, line, letter, lip, light, lunch
man, map, moon, milk, mouse, mirror, mountain, market, meal, music, magnet, match
net, nose, nest, name, nail, night, number, note, neck, nurse, noise, nation
owl, oil, oven, orange, ocean, order, orbit, open, option, owner, object, office
pen, pig, pot, plate, plane, pumpkin, pearl, park, path, piano, point, paper
queen, quill, quiz, quilt, quiet, quick, quote, quest, queue, quake, quart, quark
rat, ring, rain, river, rope, road, rose, rock, rule, room, root, radio
sun, sock, star, ship, shoe, stone, sugar, song, salt, sand, seed, snake
top, toy, tree, train, table, tube, tiger, tool, time, tent, team, towel
urn, use, unit, uncle, under, upper, uniform, union, urban, urge, ultra, usual
van, vase, veil, voice, valley, visit, value, vest, vote, view, vine, victory
wax, web, wall, wind, water, wheel, wave, wolf, wing, worm, word, wood
xylophone, xerox
yam, yard, yarn, yawn, year, yellow, yogurt, yolk, youth, yield, yeti, yoga
zoo, zebra, zone, zero, zip, zinc, zeal, zest, zigzag, zoom, zombie, zodiac
C.4 Additional results and plots for in-context prompting (Section 4.1)
C.4.1 Additional results: character-level vocabulary
Figure 10: In-context accuracy for Qwen2.5‑7B and Llama‑3-8B across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 11: In context accuracy for Qwen2.5‑7B-Instruct and Llama‑3-8B-Instruct across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 12: In context accuracy for Qwen2.5‑32B-Instruct and Llama‑3-70B-Instruct across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
As evident from all the figures here (smaller completion models - Figure 10), (smaller instruct models - Figure 11), (bigger instruct models - Figure 12), the Directional Bias persists across models sizes and types. While in the smaller models the gaps smaller for the retrieval tasks, nevertheless there is still a pronounced gap within the copying task variants.
Amongst the prompts there was high variability for all our retrieval setups, but close to none for the copying setups. Our bare template performed the worst in most retrieval setups, and the highest accuracy metrics were achieved when the instruction template was – MATH_RULE_EXPLAINED, with the few shot setting as SMALL with the presence of a separator. For copying, just the BARE setup was enough to get high accuracy numbers without the need to specify any instructions (especially for the forward copying cases). Providing instructions did improve the performance especially for the Qwen2.5-Instruct models.
We also carried out the same experiments with newer Qwen3 [Team, 2025] models with “thinking” variants to see if our findings still hold. The results from Figures 13, 14, 15, 16 show that even the model trained to explicitly generate a chain-of-thought for solving task still have the same Directional and Uniqueness biases.
Figure 13: In-context accuracy for Qwen3‑14B across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 14: In-context accuracy for Qwen3‑4B across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 15: In context accuracy for Qwen3‑4B-Thinking across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 16: In context accuracy for Qwen3‑14B-Thinking across all our character-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
C.4.2 Results: word-level vocabulary
We find that the core asymmetries persist even in the more naturalistic word-level setting (see Figures 17, 18, 19, 20, 21, 22, 23, 24). Although length generalization improves slightly for the larger models, the pretraining induced Directional Bias and the architecture-driven Uniqueness Bias remain clearly evident across all LLMs.
Figure 17: In context accuracy for Qwen2.5‑32B and Llama‑3-70B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 18: In-context accuracy for Qwen2.5‑7B and Llama‑3-8B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 19: In context accuracy for Qwen2.5‑7B-Instruct and Llama‑3-8B-Instruct across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 20: In context accuracy for Qwen2.5‑32B-Instruct and Llama‑3-70B-Instruct across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 21: In-context accuracy for Qwen3‑14B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 22: In-context accuracy for Qwen3‑4B across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 23: In context accuracy for Qwen3‑4B-Thinking across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
Figure 24: In context accuracy for Qwen3‑14B-Thinking across all our word-level tasks averaged over 3 seeds and prompt variations per task. The same patterns as discussed in Section 4.1 are visible.
C.5 Promping: Lorem Ipsum Copying (Section 4.2)